Use x and y to solve the following problem. Solve with a system of two equations. A candy mix sells for $1.10 per pound. If the mix is composed of two kinds of candy, one worth $.90 per pound and the other worth $1.50 per pound, how many pounds of each would be in a 60 pound mixture?
Answer by mananth(16946)

candy I = x lb
candy II= y lb
x+y=60
0.9x+1.5y=60*1.10

1 x + 1 y = 60 .............1
0.9 x + 1.5 y = 66 .............2

multiply (1)by -1.5
Multiply (2) by 1
-1.5 x -1.5 y = -90
0.9 x 1.5 y = 66
Add the two equations
-0.6 x = -24
/ -0.6
x = 40 lb
plug value of x in (1)
1 x + 1 y = 60
40 + 1 y = 60
1 y = 60 -40
1 y = 20
y = 20 lb